Mr Charles borrowed the sum of $10000 from a finance company for 6 years and 8% per annum at simple interest rate. The amount returned by him to the finance company is
Answer by math_tutor2020(3817)


Use the simple interest formula
i = P*r*t
i = 10,000*0.08*6
i = 4,800

Mr Charles has to pay $4,800 in simple interest.
This is on top of the principal $10,000 he borrowed.

In total, he pays back 10,000+4,800 = $14,800

Another slightly alternative formula to use is this:
A = P*(1+r*t)
A = 10,000*(1+0.08*6)
A = 14,800